# A Survey on Quantum Machine Learning: Basics, Current Trends, Challenges, Opportunities, and the Road Ahead

### v2

@@ Abstract
Index Terms:
Quantum Computing, Quantum Computer, Quantum Machine Learning, Quantum Neural Networks, Machine Learning, Neural Networks, Quantum Supremacy, Qubit, Superposition, Quantum Correlation, Entanglement, Quantum Gate, Quantum Circuit, Quantum Noise, Noisy Intermediate-Scale Quantum, Fault-Tolerant Quantum Computing, Parametrized Quantum Circuits, Quantum Annealing, Quantum Kernels, Variational Quantum Eigensolver, Quantum Data, Quantum Encoding, Quantum Datasets, Quantum Hardware, Quantum Simulator, Qiskit, PennyLane, Quantum Applications.
I Introduction
Machine Learning (ML) systems are well-established tools for identifying patterns in data and generalizing complex, nonlinear problems. These systems have found applications in various domains, including computer vision, healthcare, finance, and the automotive industry. However, ML practitioners must navigate the substantial computing resources required to run large ML models. Despite employing numerous hardware-aware optimizations such as compression and approximations [1, 2, 3, 4, 5], the limitations of current computing infrastructures and technologies constrain the computational capabilities of ML systems.
The high computational demands of modern ML models necessitate advanced hardware development. According to Moore’s law, the number of transistors on an integrated circuit doubles approximately every two years. However, this trend has reached its limits with current high-end CPUs and GPUs [6]. This physical saturation restricts computational power, leading to delays in processing, developmental, and scientific discovery within the ML community. For instance, training Large Language Models with hundreds of billions of parameters and trillions of tokens is extremely compute-intensive [7]. Such tasks require massive investments in time and hardware resources, which only a few high-end companies can afford. To overcome these physical limits and support further discoveries, there is an urgent need to explore new technological avenues of hardware systems that can enhance the computational efficiency for solving given real-world problems by closely simulating and comprehending them.
TABLE I: Qualitative comparison between our work and related QML surveys.
@@ I Introduction
[8] 2006 12 24
[9] 2014 19 71
[10] 2015 38 97
[11] 2016 17 100
[12] 2017 106 309
[13] 2018 33 187
[14] 2019 15 74
[15] 2019 29 309
[16] 2020 8 71
[17] 2020 6 47
[18] 2020 17 165
[19] 2021 9 75
[20] 2021 4 41
[21] 2021 14 73
[22] 2022 6 21
[23] 2022 17 154
[24] 2022 12 59
[25] 2023 19 116
[26] 2023 49 199
[27] 2023 337 851
[28] 2023 59 301
[29] 2023 21 87
[30] 2024 44 42
Our work 2024 35 187
One of the most promising solutions to this bottleneck is Quantum Computing (QC). Initially proposed by Feynman [31] and further developed by Preskill [32, 33], QC systems exploit quantum mechanical phenomena to significantly improve performance and information processing compared to classical systems. Unlike classical systems that struggle to encompass natural processes, quantum computers operate on similar principles to those found in nature. This similarity suggests that QC could help us better understand natural phenomena. Moreover, QC has the potential to solve problems more cost-efficiently through optimizations, time savings, and environmentally friendly designs.
The Quantum Machine Learning (QML) paradigm represents an excellent opportunity for researchers and industries to achieve remarkable discoveries and design efficient solutions for complex real-world problems. QML systems, driven towards practicality and improved performance over classical systems, open new avenues for the community to discover, build, and align their designs across different levels of the quantum stack [34]. This integration of QC and ML could lead to groundbreaking advancements and a deeper understanding of the world around us.
I-A Scope and Contributions
@@
The primary goal of this survey is to introduce and provide substantial knowledge on QML, laying a solid foundation for solving new and advanced problems. In contrast to prior QML surveys [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30], we provide a well-structured and thoroughly comprehensive collection of information and concepts that summarize the current state of development of QML. Table I illustrates the comparison between this work and a selection of related QML surveys, highlighting the features and topics each covers. It becomes evident from the table that previous surveys have at least one missing feature, whereas this work delivers a detailed review encompassing all key aspects of QML. The contributions of this paper are summarized below.
@@
I-B Outline
This paper is organized into different sections and sub-sections. Section I introduces the problem, the scope, and the contributions of this paper. Section II presents foundational theoretical concepts of general quantum computers, the challenges of the current technologies, and the state-of-the-art methodologies to overcome these issues. Section III provides an overview of the claimed and proven advantages of QC compared to classical computing. This section discusses the motivations that are driving researchers and industries to design QC and QML systems. Section IV discusses the plethora of the most common QML algorithms that are present in the literature and provides an overview of their applicability. Section V provides the necessary information to understand the existing methods for manipulating quantum data in a way that it can be used in the QML workflow, along with a curated list of dataset resources from the QML perspective. Section VI introduces the existing tools and technologies that are available for QML practitioners to experiment with and investigate further. Section VII presents the current and potential applications of QML that majorly highlight their benefits over classical computing. Section VIII concludes the paper and provides future outlooks in the QML field.
II Quantum Computing Preliminaries
@@
II-A Understanding the Qubit
@@
On the other hand, the resultant state of the quantum computation system’s processing is dictated by the probabilistic occurrence of the collective system’s states as a solution over multiple experiment sample runs (called shots). For each shot, the resulting solution over the super-positioned states, at the final computational stage of the system circuit after state manipulation (quantum processing) using a set of gates, passes through the measurement operation to form the measured state. It is represented through the basis states |$\ket{0}$ and |$\ket{1}$, followed by the final probabilistic distribution over possible measured state outcomes as solutions to the problem [35].
Figure 1: (a) Classical bit vs. (b) Quantum Bit with $\alpha$ and $\beta$ as respective amplitude values of a state-vector, creating a superposition state where $|\alpha|^{2}+|\beta|^{2}=1$ as per the Max-Born Rule, satisfying the completeness equation which states to have all probabilities summing up to 1 [36]. $|\alpha|^{2}$ and $|\beta|^{2}$ give the respective probability for each state.
@@
Figure 2: Bloch sphere for a single two-level qubit system.
II-B Superposition
@@
The Wave-Particle duality theory posits that waves can exhibit particle-like properties and particles can exhibit wave-like properties. This duality is a cornerstone of quantum mechanics, contrasting sharply with the principles of classical mechanics or Newtonian physics [37]. In classical mechanics, waves and particles are distinct entities with no overlap in their properties. Quantum mechanics, however, reveals that entities such as photons and electrons can behave both as particles and as waves, depending on the experimental setup. This dual nature is exemplified in phenomena such as the double-slit experiment, where particles like electrons create an interference pattern characteristic of waves when not observed, but act like particles when observed.
The Max-Born rule provides a probabilistic framework for quantum mechanics. It states that the probability density of finding a quantum system in a given state is proportional to the square of the amplitude of the system’s wave function at that state. This is usually represented within a state-vector formalism in Hilbert space. The rule is fundamental to the measurement process in quantum mechanics [38]. When a qubit is measured, the probability of finding it in a particular state is determined by the square of the amplitude of its wave function for that state.
@@
II-B1 Single Qubit Superposition
@@ II-B2 Two-Qubit Superposition
II-C Quantum Gates
Quantum logic gates are one of the essential parts of a quantum computer and are the building blocks of all quantum algorithms. These gates are mathematically described by unitary matrices, and their action is always logically reversible. The gates correspond to operators used to manipulate the quantum state of qubits. The only condition on a valid gate (and its corresponding quantum operator) is that it should be unitary. By unitary we mean that it meets the $U\textsuperscript{\textdagger}=U^{-1}$ condition, where $U$ is the gate operation matrix and $U\textsuperscript{\textdagger}$ is the adjoint of $U$. This property also ensures that it is a reversible gate operation [39].
II-D Quantum Circuits
A quantum circuit is a collection of quantum gates interconnected by quantum wires. The actual structure of a quantum circuit, the number and the types of gates, as well as the interconnection scheme, are dictated by the unitary transformation, U, executed by the circuit [40]. A circuit is an independent module composed of gates, wires, and qubits arranged in a certain way to perform a task. Hence, its functionality is similar to what we refer to as an algorithm. A collection of circuits are independent functional quantum units using the output of a circuit as the input of another one to form a complete algorithm for the solution of the given problem [41] (see Figure 7).
@@ II-D Quantum Circuits
II-E Quantum Correlations
@@
In contrast, quantum mechanics introduces a probabilistic nature to correlation. Observables in a quantum system do not have a single, definite outcome [42]. Instead, they have a range of possible values, known as eigenvalues, which are associated with certain probabilities. These eigenvalues are the result of the system being represented by a Hermitian matrix. Thus, even without any perturbation, measurements of a quantum system can yield different results over time. These varying outcomes are described by eigenfunctions that exist within a Hilbert space, a mathematical framework that encompasses all possible states of the system.
II-E1 Quantum Entanglement
Quantum entanglement is a central concept in quantum information theory and represents one of the most counterintuitive aspects of quantum mechanics. Entanglement occurs when two particles become linked such that the state of one particle instantaneously influences the state of the other, regardless of the distance separating them [43, 44, 45]. This phenomenon, which Albert Einstein famously referred to as a “spooky action at a distance”, defies the classical notion of local realism [46].
The mechanism of quantum entanglement is illustrated in Figure 8, showing how changes in one particle affect its entangled partner. While entanglement itself is nondeterministic, meaning the specific outcomes cannot be predicted with certainty, achieving deterministic control over entangled states could lead to groundbreaking applications across various fields. Quantum development tools like Qiskit [47] and PennyLane [48] enable researchers to create and manipulate entangled states for solving real-world problems.
@@ II-E1 Quantum Entanglement
II-E2 Quantum Decoherence
@@
Quantum decoherence is the process by which a quantum system loses its coherence due to interactions with the external environment. This loss of coherence manifests as the system’s transition from a pure quantum state to a mixed state, effectively causing a collapse of the wave function [49]. Decoherence can be viewed as the leakage of information from the quantum system into the environment, or as environmental interference disrupting the quantum state.
@@
II-F Quantum Noise
Leading from the idea of coherence and decoherence, quantum noise makes the implementation of large-scale and reliable quantum computers with low error rates extremely challenging. Therefore, it is imperative to properly characterize the noise in quantum systems and devise mitigation techniques to detect and correct the errors [50].
II-F1 Uncertainty Principle
@@
Heisenberg’s uncertainty principle states that it is fundamentally impossible to simultaneously measure the exact value of certain pairs of related physical quantities, such as position ($x$) and momentum ($p$), with infinite precision [51]. Mathematically, this principle is expressed as in Equation 2:
@@
II-F2 Noisy Intermediate-Scale Quantum Era
The long-term goal for QC is to develop functional quantum algorithms that can solve problems despite the noisy environmental systems they work in. Completely eliminating the noise errors is extremely difficult. Hence, the community has defined a set of achievable intermediate goals to evolve from Noisy Intermediate-Scale Quantum (NISQ) [^1] NISQ is a term coined by John Preskill [32]. to Fault Tolerant Quantum Computers (FTQC); see Figure 9.
The current state of quantum computing is referred to as the NISQ era [32]. Current quantum processors only support $50$-$100$’s of qubits but are not advanced enough to guarantee complete fault tolerance. Despite that, they represent a valid infrastructure for experimenting and improving the designs towards ideal systems. In the NISQ era, near-term hybrid quantum-classical algorithms are designed and applied to various fields, such as quantum chemistry, QML, and combinatorial optimization.
@@ II-F2 Noisy Intermediate-Scale Quantum Era
II-F3 Error Correction
@@
Due to the no-cloning theorem [52], creating a perfect copy of a quantum state is impossible. Hence, Quantum Error Correction (QEC) codes introduce redundancy by spreading the information of a single qubit onto an entangled state of multiple qubits. With this approach, it is possible to perform multi-qubit (syndrome) measurements to extract the information about the error without altering the quantum information. Typical QEC schemes are the three-qubit bit flip code [53], the three-qubit phase flip code [54] and Shor’s nine qubit code [55].
@@
Figure 10: Current state of QEC development. The figure shows the physical error rate threshold per gate that is currently not possible to overcome in the NISQ era.
II-F4 Error Mitigation
Despite sharing the same goal as QEC to reduce the impact of quantum noise, Quantum Error Mitigation (QEM) techniques operate differently. While QEC aims to restore the correct value after the error occurs, QEM aims to reduce or suppress the errors that occur during computation. Moreover, as shown in Figure 11, QEC completely removes the noise, while QEM keeps a small amount of noise under control. A scenario where QEC is applied, called ‘hard regime’, is theoretically ideal but difficult to realize. The other scenario where QEM is employed, called ‘easy regime’, is experimentally easier to implement but at the price of lower qubit protection. For QEM deployment, Mitiq [56] is an open-source error mitigation package in Python that implements QEM techniques like zero-noise extrapolation [57] and probabilistic error cancellation [58] for quantum computers.

### v3

@@ Abstract
1 Introduction
Machine Learning (ML) systems are well-established tools for identifying patterns in data and generalizing complex, nonlinear problems. These systems have found applications in various domains, including computer vision, healthcare, finance, and the automotive industry. However, ML practitioners must navigate the substantial computing resources required to run large ML models. Despite employing numerous hardware-aware optimizations such as compression and approximations (Marchisio et al., 2018; Hanif et al., 2018; Marchisio et al., 2019; Hanif et al., 2022; Leon et al., 2023), the limitations of current computing infrastructures and technologies constrain the computational capabilities of ML systems.
The high computational demands of modern ML models necessitate advanced hardware development. According to Moore’s law, the number of transistors on an integrated circuit doubles approximately every two years. However, this trend has reached its limits with current high-end CPUs and GPUs (Gustafson, 2011). This physical saturation restricts computational power, leading to delays in processing, developmental, and scientific discovery within the ML community. For instance, training Large Language Models with hundreds of billions of parameters and trillions of tokens is extremely compute-intensive (OpenAI, 2023). Such tasks require massive investments in time and hardware resources, which only a few high-end companies can afford. To overcome these physical limits and support further discoveries, there is an urgent need to explore new technological avenues of hardware systems that can enhance the computational efficiency for solving given real-world problems by closely simulating and comprehending them.
Table 1: Qualitative comparison between our work and related QML surveys.
@@ 1 Introduction
(Aïmeur et al., 2006) 2006 12 24
(Schuld et al., 2014) 2014 19 71
(Adcock et al., 2015) 2015 38 97
(Montanaro, 2016) 2016 17 100
(Dunjko & Briegel, 2017) 2017 106 309
(Ciliberto et al., 2018) 2018 33 187
(Jeswal & Chakraverty, 2018) 2019 15 74
(Resch & Karpuzcu, 2019) 2019 29 309
(Ramezani et al., 2020) 2020 8 71
(Chakraborty et al., 2020) 2020 6 47
(Bharti et al., 2020) 2020 17 165
(Mangini et al., 2021) 2021 9 75
(Kwak et al., 2021) 2021 4 41
(Zhao & Wang, 2021) 2021 14 73
(N et al., 2022) 2022 6 21
(Umer & Sharif, 2022) 2022 17 154
(Schetakis et al., 2022) 2022 12 59
(Markidis, 2023) 2023 19 116
(Massoli et al., 2023) 2023 49 199
(Dalzell et al., 2023) 2023 337 851
(Huynh et al., 2023) 2023 59 301
(Tychola et al., 2023) 2023 21 87
(Evans et al., 2024) 2024 44 42
(Wang & Liu, 2024) 2024 30 165
(Chen et al., 2024) 2024 44 134
Our work 2025 54 191
One of the most promising solutions to this bottleneck is Quantum Computing (QC). Initially proposed by Feynmann (1982) and further developed by Preskill (2018; 2023), QC systems exploit quantum mechanical phenomena to significantly improve performance and information processing compared to classical systems. Unlike classical systems that struggle to encompass natural processes, quantum computers operate on similar principles to those found in nature. This similarity suggests that QC could help us better understand natural phenomena. Moreover, QC has the potential to solve problems more cost-efficiently through optimizations, time savings, and environmentally friendly designs.
The Quantum Machine Learning (QML) paradigm represents an excellent opportunity for researchers and industries to achieve remarkable discoveries and design efficient solutions for complex real-world problems. QML systems, driven towards practicality and improved performance over classical systems, open new avenues for the community to discover, build, and align their designs across different levels of the quantum stack (Schuld & Killoran, 2022). This integration of QC and ML could lead to groundbreaking advancements and a deeper understanding of the world around us.
1.1 Scope and Contributions
@@
The primary goal of this survey is to introduce and provide substantial knowledge on QML, laying a solid foundation for solving new and advanced problems. In contrast to prior QML surveys (Aïmeur et al., 2006; Schuld et al., 2014; Adcock et al., 2015; Montanaro, 2016; Dunjko & Briegel, 2017; Ciliberto et al., 2018; Jeswal & Chakraverty, 2018; Resch & Karpuzcu, 2019; Ramezani et al., 2020; Chakraborty et al., 2020; Bharti et al., 2020; Mangini et al., 2021; Kwak et al., 2021; Zhao & Wang, 2021; N et al., 2022; Umer & Sharif, 2022; Schetakis et al., 2022; Markidis, 2023; Massoli et al., 2023; Dalzell et al., 2023; Huynh et al., 2023; Tychola et al., 2023; Evans et al., 2024; Wang & Liu, 2024; Chen et al., 2024), we provide a well-structured and thoroughly comprehensive collection of information and concepts that summarize the current state of development of QML. Table 1 illustrates the comparison between this work and a selection of related QML surveys, highlighting the features and topics each covers. It becomes evident from the table that previous surveys have at least one missing feature, whereas this work delivers a detailed review encompassing all key aspects of QML. The contributions of this paper are summarized below.
@@
1.2 Outline
This paper is organized into different sections and sub-sections. Section 1 introduces the problem, the scope, and the contributions of this paper. Section 2 presents foundational theoretical concepts of general quantum computers, the challenges of the current technologies, and the state-of-the-art methodologies to overcome these issues. Section 3 provides an overview of the claimed and proven advantages of QC compared to classical computing. This section discusses the motivations that are driving researchers and industries to design QC and QML systems. Section 4 discusses the plethora of the most common QML algorithms that are present in the literature and provides an overview of their applicability. Section 5 provides the necessary information to understand the existing methods for manipulating quantum data in a way that it can be used in the QML workflow, along with a curated list of dataset resources from the QML perspective. Section 6 introduces the existing tools and technologies that are available for QML practitioners to experiment with and investigate further. Section 7 presents the current and potential applications of QML that majorly highlight their benefits over classical computing. Section 8 concludes the paper and provides future outlooks in the QML field.
2 Quantum Computing Preliminaries
@@
2.1 Understanding the Qubit
Figure 1: (a) Classical bit vs. (b) Quantum Bit with $\alpha$ and $\beta$ as respective amplitude values of a state-vector, creating a superposition state where $|\alpha|^{2}+|\beta|^{2}=1$ as per the Max-Born Rule, satisfying the completeness equation which states to have all probabilities summing up to 1 (Nielsen & Chuang, 2010b). $|\alpha|^{2}$ and $|\beta|^{2}$ give the respective probability for each state.
@@ 2.1 Understanding the Qubit
On the other hand, the resultant state of the quantum computation system’s processing is dictated by the probabilistic occurrence of the collective system’s states as a solution over multiple experiment sample runs (called shots). For each shot, the resulting solution over the super-positioned states, at the final computational stage of the system circuit after state manipulation (quantum processing) using a set of gates, passes through the measurement operation to form the measured state. It is represented through the basis states |$\ket{0}$ and |$\ket{1}$, followed by the final probabilistic distribution over possible measured state outcomes as solutions to the problem (Marinescu & Marinescu, 2012a).
Figure 2: Bloch sphere for a single two-level qubit system.
@@
2.2 Superposition
@@
The Wave-Particle duality theory posits that waves can exhibit particle-like properties and particles can exhibit wave-like properties. This duality is a cornerstone of quantum mechanics, contrasting sharply with the principles of classical mechanics or Newtonian physics (Wheaton, 2009). In classical mechanics, waves and particles are distinct entities with no overlap in their properties. Quantum mechanics, however, reveals that entities such as photons and electrons can behave both as particles and as waves, depending on the experimental setup. This dual nature is exemplified in phenomena such as the double-slit experiment, where particles like electrons create an interference pattern characteristic of waves when not observed, but act like particles when observed.
The Max-Born rule provides a probabilistic framework for quantum mechanics. It states that the probability density of finding a quantum system in a given state is proportional to the square of the amplitude of the system’s wave function at that state. This is usually represented within a state-vector formalism in Hilbert space. The rule is fundamental to the measurement process in quantum mechanics (Stoica, 2023). When a qubit is measured, the probability of finding it in a particular state is determined by the square of the amplitude of its wave function for that state.
@@
2.2.1 Single Qubit Superposition
@@ 2.2.2 Two-Qubit Superposition
2.3 Quantum Gates
Quantum logic gates are one of the essential parts of a quantum computer and are the building blocks of all quantum algorithms. These gates are mathematically described by unitary matrices, and their action is always logically reversible. The gates correspond to operators used to manipulate the quantum state of qubits. The only condition on a valid gate (and its corresponding quantum operator) is that it should be unitary. By unitary we mean that it meets the $U\textsuperscript{\textdagger}=U^{-1}$ condition, where $U$ is the gate operation matrix and $U\textsuperscript{\textdagger}$ is the adjoint of $U$. This property also ensures that it is a reversible gate operation (Barenco et al., 1995).
2.4 Quantum Circuits
@@
A quantum circuit is a collection of quantum gates interconnected by quantum wires. The actual structure of a quantum circuit, the number and the types of gates, as well as the interconnection scheme, are dictated by the unitary transformation, U, executed by the circuit (Marinescu & Marinescu, 2012b). A circuit is an independent module composed of gates, wires, and qubits arranged in a certain way to perform a task. Hence, its functionality is similar to what we refer to as an algorithm. A collection of circuits are independent functional quantum units using the output of a circuit as the input of another one to form a complete algorithm for the solution of the given problem (Muñoz Coreas & Thapliyal, 2022) (see Figure 7).
2.5 Quantum Correlations
@@
In contrast, quantum mechanics introduces a probabilistic nature to correlation. Observables in a quantum system do not have a single, definite outcome (Adesso et al., 2016). Instead, they have a range of possible values, known as eigenvalues, which are associated with certain probabilities. These eigenvalues are the result of the system being represented by a Hermitian matrix. Thus, even without any perturbation, measurements of a quantum system can yield different results over time. These varying outcomes are described by eigenfunctions that exist within a Hilbert space, a mathematical framework that encompasses all possible states of the system.
2.5.1 Quantum Entanglement
Quantum entanglement is a central concept in quantum information theory and represents one of the most counterintuitive aspects of quantum mechanics. Entanglement occurs when two particles become linked such that the state of one particle instantaneously influences the state of the other, regardless of the distance separating them (Bussandri & Lamberti, 2021; Cacciapuoti et al., 2020; Illiano et al., 2022). This phenomenon, which Albert Einstein famously referred to as a “spooky action at a distance”, defies the classical notion of local realism (Einstein et al., 1971).
@@ 2.5.1 Quantum Entanglement
The mechanism of quantum entanglement is illustrated in Figure 8, showing how changes in one particle affect its entangled partner. While entanglement itself is nondeterministic, meaning the specific outcomes cannot be predicted with certainty, achieving deterministic control over entangled states could lead to groundbreaking applications across various fields. Quantum development tools like Qiskit (Qiskit contributors, 2023) and PennyLane (Bergholm et al., 2018) enable researchers to create and manipulate entangled states for solving real-world problems.
2.5.2 Quantum Decoherence
@@
Quantum decoherence is the process by which a quantum system loses its coherence due to interactions with the external environment. This loss of coherence manifests as the system’s transition from a pure quantum state to a mixed state, effectively causing a collapse of the wave function (Schlosshauer, 2019). Decoherence can be viewed as the leakage of information from the quantum system into the environment, or as environmental interference disrupting the quantum state.
@@
2.6 Quantum Noise
Leading from the idea of coherence and decoherence, quantum noise makes the implementation of large-scale and reliable quantum computers with low error rates extremely challenging. Therefore, it is imperative to properly characterize the noise in quantum systems and devise mitigation techniques to detect and correct the errors (Shaib et al., 2021).
2.6.1 Uncertainty Principle
@@
Heisenberg’s uncertainty principle states that it is fundamentally impossible to simultaneously measure the exact value of certain pairs of related physical quantities, such as position ($x$) and momentum ($p$), with infinite precision (Heisenberg, 1925). Mathematically, this principle is expressed as in Equation 2:
@@
2.6.2 Noisy Intermediate-Scale Quantum Era
@@
The long-term goal for QC is to develop functional quantum algorithms that can solve problems despite the noisy environmental systems they work in. Completely eliminating the noise errors is extremely difficult. Hence, the community has defined a set of achievable intermediate goals to evolve from Noisy Intermediate-Scale Quantum (NISQ) [^1] NISQ is a term coined by John Preskill (Preskill, 2018). to Fault Tolerant Quantum Computers (FTQC); see Figure 9.
The current state of quantum computing is referred to as the NISQ era (Preskill, 2018). Current quantum processors only support $50$-$100$’s of qubits but are not advanced enough to guarantee complete fault tolerance. Despite that, they represent a valid infrastructure for experimenting and improving the designs towards ideal systems. In the NISQ era, near-term hybrid quantum-classical algorithms are designed and applied to various fields, such as quantum chemistry, QML, and combinatorial optimization.
2.6.3 Error Correction
@@
Figure 10: Current state of QEC development. The figure shows the physical error rate threshold per gate is currently not possible to overcome in the NISQ era.
Due to the no-cloning theorem (Wootters & Zurek, 1982), creating a perfect copy of a quantum state is impossible. Hence, Quantum Error Correction (QEC) codes introduce redundancy by spreading the information of a single qubit onto an entangled state of multiple qubits. With this approach, it is possible to perform multi-qubit (syndrome) measurements to extract the information about the error without altering the quantum information. Typical QEC schemes are the three-qubit bit flip code (Peres, 1985), the three-qubit phase flip code (Nielsen & Chuang, 2010a) and Shor’s nine qubit code (Shor, 1995).
@@
2.6.4 Error Mitigation